Fuji apples cost $3.00 per pound, and Golden Delicious apples cost $2.00 per pound. A childcare center purchases 30 pounds of a combination of the two types of apples for a total of $80.

Which value could replace x in the table?

30  g + g

3(30  g)

30  g(3)

30  + 3

1
Expert's answer
2021-01-29T05:08:34-0500

Solution-

Given Fuji apple cost $3 per pound

Golden Apple cost $2 per pound

Purchase 30 pound in $80

Let quantity of fuji Apple is f

and golden Apple is g

So Equation's

f+g=30.............(1)

3f+2g=80...........(2)

After solve this equation we get

f=20 and g=10

Put value of f in equation (1) and multiple by 3

New equation is

3×20+3(g)=90

=90-60-3(g)

=30-3(g)

This 30-3(g) value can put in the table.